Can you show me how to do the following:
all of this squared (x^2+3x+6) = 4
Don't you sq what is in the paren. and then sq 4 also to solve?

Answer by sabanasir(22)

absolutely u shall do it the way u say so only.
as i understand u mean to say the question is:
[(x^2+3x+6) = 4 ] ^2 ???
and you say we square root on either sides??
yes thats wat we shall do.
then we get:
x^2 + 3x + 6 = 4
x^2 + 3x = 4 -6
x^2 + 3x = -2
x^2 + 3x + 2 = 0
x^2 + x + 2x + 2 = 0
x(x+1) + 2(x+1)=0
(x+1)(x+2)=0
x=-1, x=-2